OpenStudy (anonymous):

please help write a direct or inverse variation equation for the set of numbers x= 3, 4, 8 y= 24,18, 9

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

never mind i think i figured it out.

OpenStudy (whpalmer4):

If you have an inverse variation, \[y = k/x\] for some value of \(k\). You can find \(k\) by plugging in corresponding values of \(x\) and \(y\): \[x=3, y = 24\]\[24 = k/3\]\[k=72\] Try out with the others: \[y=72/x\]\[18=72/4\checkmark\]\[9=72/8\checkmark\]